### v2

# A System-of-Systems Convergence Paradigm for Societal Challenges of the Anthropocene

Abstract
Modern societal challenges, such as climate change, urbanization, and water resource management, demand integrated, multi-discipline, multi-problem approaches to frame and address their complexity. Unfortunately, current methodologies often operate within disciplinary silos, leading to fragmented insights and missed opportunities for convergence. A critical barrier to cross-disciplinary integration lies in the disparate ontologies that shape how different fields conceptualize and communicate knowledge. To address these limitations, this paper proposes a system-of-systems (SoS) convergence paradigm grounded in a meta-cognition map, a framework that integrates five complementary domains: real-world observations, systems thinking, visual modeling, mathematics, and computing. The paradigm is based on the Systems Modeling Language (SysML), offering a standardized, domain-neutral approach for representing and analyzing complex systems. The proposed methodology is demonstrated through a case study of the Chesapeake Bay Watershed, a socio-environmental system requiring coordination across land use, hydrology, economic and policy domains. By modeling this system with SysML, the study illustrates practical strategies for navigating interdisciplinary challenges and highlights the potential of agile SoS modeling to support large-scale, multi-dimensional decision-making.
[…]
Earth’s Future
Department of Civil and Environmental Engineering, Virginia Tech, Blacksburg, Virginia 24061, United States School of Systems and Enterprises, Stevens Institute of Technology, Hoboken, New Jersey, 07030, United States Askew School of Public Administration and Policy, Florida State University, Tallahassee, Florida 32306 Department of Environment and Sustainability, University at Buffalo, Buffalo, New York, 14260, United States Biodiversity and Natural Resources Program, International Institute for Applied Systems Analysis (IIASA), Schlossplatz 1, A-2361 Laxenburg, Austria Commonwealth Scientific and Industrial Research Organisation (CSIRO), Private Bag 10, Clayton South, Victoria 3169, Australia Departments of Oncology, Industrial and Systems Engineering, and Computer Science, Wayne State University, Detroit, MI 48201
[…]
Megan S. Harrismegstephh@vt.edu
{keypoints}
Fragmented ontologies hinder integration across socio environmental systems of systems
We develop a convergence framework using Systems Modeling Language and a meta cognition map
Application to the Chesapeake Bay Region demonstrates scalable interdisciplinary modeling
Plain Language Summary
Managing regional watersheds such as the Chesapeake Bay is challenging because water quality, land use, ecology, and policy decisions are tightly interconnected. Many existing modeling tools estimate nutrient loads effectively but are difficult to integrate with broader analyses of social and environmental systems. This limits their usefulness for coordinated decision making.
This study introduces a new modeling approach designed to bridge disciplinary boundaries. The framework combines systems thinking, visual modeling, mathematics, computing, and real world data within a unified system of systems structure. This structure helps researchers and decision makers represent complex watershed interactions in a more transparent and extensible way.
We demonstrate the approach using the Chesapeake Bay Watershed. Results show how the framework can support more integrated analysis of coupled human and natural processes. The method provides a foundation for future decision support tools aimed at improving large scale environmental management.
1 Introduction
The Earth has entered a new era, the Anthropocene, marked by profound and irreversible human-driven transformations of the natural and built systems of the planet [Folke:2021:00, demos:2017:00, williams:2022:00, steffen:2011:00, zalasiewicz:2017:00]. Climate change, eutrophication, food insecurity, water scarcity, and energy insecurity are not isolated crises: they are tightly interwoven and mutually reinforcing [Wang:2019:10, stavi:2021:00, ho:2019:00]. These Anthropocene challenges span tightly coupled ecological, economic, industrial, and socio-cultural systems (Fig. 1), none of which can be addressed in isolation [berkes:2017:00, sterner:2019:00, steffen:2011:00]. Instead, they demand a systems-based understanding of how diverse domains interact across space, time, and institutional contexts [Cornell:2013:00, plummer:2007:00, wu:2019:01].
Figure 1: Graphical illustration of an interconnected geophysical, biophysical, sociocultural and sociotechnical system of Anthropocene systems.
Despite efforts and progresses to develop integrated approaches across various domains [yang:2025:00, razavi:2025:00, schluter:2023:00], responses still remain mostly isolated in policy and practice. Approaches are often framed as single-discipline, single-problem interventions, supported by siloed institutions and legacy analytical tools [toivanen:2017:00, verburg:2016:00]. Even well-meaning convergence initiatives that focus on one societal challenge are frequently uncoordinated in practice and policy with efforts in adjacent challenges like water access, food security, or climate adaptation [Bi:2022:00, Farid:2022:01]. These fragmented efforts are not only inefficient, but can also be counterproductive, as progress in one domain may inadvertently undermine others [Albanese:2022:00, wright:2018:00, lim:2018:00]. For example, to cope with water scarcity, regions often rely on energy-intensive water solutions that can amplify greenhouse gas emissions and climate change.
This misalignment stems from the fact that Anthropocene systems are not only interdependent in their physical dynamics, but also in the institutions, policies, and communities that govern them (Fig. 1) [dryzek:2016:00, brondizio:2016:00]. Further, as illustrated in Fig. 2 (adapted from [weck:2011:00]), disciplinary tunnel vision obscures these interconnections, preventing us from recognizing root causes, system-wide feedbacks, and cross-sector synergies. For example, single-discipline approaches may model watershed nutrient transport, management policy, climate-driven runoff, or ecological degradation, but rarely integrate all of these processes. In a world of tightly coupled environmental and societal systems, approaches that treat problems in isolation risk missing tradeoffs, squandering opportunities for alignment, and triggering maladaptive outcomes.
To avoid these pitfalls, a convergence paradigm must evolve from single-discipline-single-problem thinking, to multi-discipline-single-problem strategies, and ultimately to multi-discipline-multi-problem frameworks that are capable of generating coordinated, cross-sector solutions [Cornell:2013:00, Farid:2022:01].
Figure 2: Tunnel vision when addressing Anthropocene societal challenges
1.1 Original Contribution
This paper presents a novel, generalizable system-of-systems (SoS) convergence paradigm for modeling and managing complex systems of Anthropocene systems. Grounded in a meta-cognition map that organizes knowledge generation across five domains–Real-World, Systems Thinking, Visual, Mathematics, and Computing–the paradigm provides a structured foundation for integrating disparate disciplinary insights. It introduces a cohesive modeling toolbelt comprising Network Science, Data-Driven Artificial Intelligence, Model-Based Systems Engineering (MBSE), and Hetero-functional Graph Theory (HFGT), each selected for their high convergence potential. The paradigm is demonstrated through a case study of the Chesapeake Bay Watershed, where it is used to integrate land use, watershed, and economic models within a shared ontological framework based on SysML and HFGT. Complemented by a transdisciplinary educational strategy and Team Science infrastructure, this work advances a replicable approach for training Anthropocene System Integrators and for integrating coherent, actionable knowledge across scientific, technical, and institutional domains.
1.2 Paper Outline
Section 1 introduces the motivating challenge: addressing Anthropocene problems requires moving beyond single-discipline, single-problem thinking toward integrated, SoS approaches. Section 2 identifies five common pitfalls that arise when attempting this transition, such as conceptual silos, incompatible ontologies, and mismatched modeling tools that hinder meaningful integration across disciplines. Section 3 proposes a structured convergence paradigm grounded in a meta-cognition map, which organizes knowledge production across five domains: Real-World, Systems Thinking, Visual, Mathematics, and Computing. This scaffold is used to articulate the characteristics of a successful SoS paradigm and to evaluate modeling methods with high convergence potential and describe characteristics of future pioneers of the framework. Section 4 demonstrates how this paradigm can be applied in practice through an integrated case study of the Chesapeake Bay Watershed, showcasing an SoS computational framework, decision-support system, and educational infrastructure designed to train Anthropocene System Integrators. Finally, Section 5 offers reflections on the broader implications of this work and outlines next steps for expanding, institutionalizing, and scaling the SoS convergence paradigm.
2 Common Pitfalls to Convergence
Despite growing recognition of the need for systemic approaches, most efforts to address the challenges of the Anthropocene fall short [norstrom:2014:00]. Stakeholders, ranging from communities to policy-makers, are hindered by tools and frameworks that are either too narrow in scope or too fragmented to reflect the true complexity of socio-environmental systems [Bi:2022:00, Farid:2022:01, dentoni:2018:00]. These limitations are not merely technical; they reflect deeper epistemic, institutional, and infrastructural gaps [razavi:2025:00]. Based on past research and the collective experiences of convergence science initiatives, five systemic pitfalls have been identified.
Pitfall 1: Regional and institutional context dependency. Tools and models developed for one region seldom translate to another [yates:2018:00]. Trade-offs and synergies manifest differently depending on local ecological conditions, cultural values, and institutional structures [Little:2019:00, Biermann:2022:00]. For example, the nutrient crediting framework of the Chesapeake Bay succeeds under strong interstate governance and total maximum daily load (TMDL) mandates [ferris:2024:00, stephenson:2006:00], but would fail to translate to the Mississippi Basin, where agricultural regulation and coordination are weaker [secchi:2019:00, marshall:2018:00]. Likewise, optimization models for renewable power grids are often developed within specific infrastructural and market contexts, such as the U.S. grid with centralized dispatch and formal electricity markets, and therefore cannot be directly transferred to regions with fundamentally different institutional and network structures without substantial reformulation [bhattacharyya:2021:00, oyewo:2023:00, babayomi:2023:00]. As noted at the Asia-Pacific Economic Cooperation (APEC) summit in 2022: “You can’t have one-nation solutions to issues which are global” [Albanese:2022:00]. Without deliberate international cooperation and regional adaptation, convergence science efforts risk being parochial or invalid across contexts [coccia:2016:00].
Pitfall 2: Fragmented ontologies and sectoral silos. Most tools are developed within individual sectors–electricity, water, transportation, food–without recognizing the interconnected systems they influence or depend upon [Farid:2016:ISC-BC06, Farid:2022:ISC-J49, Farid:2022:01]. This fragmentation leads to ontological mismatches when attempting integration [monteiro:2009:01]. Researchers attempting multi-sector modeling must reconcile conflicting problems, definitions, data structures, and modeling assumptions across disciplines [iyer:2019:00, reed:2022:01, srikrishnan:2022:00]. For instance, national decarbonization scenarios often integrate detailed power sector models with coarse or static representations of water and land systems, leading to unrealistic assumptions about cooling water availability or biofuel feedstock supply [kraucunas:2015:00]. Even ambitious efforts like food-energy-water nexus studies often fall short: a systematic review of 245 publications revealed that most did not actually capture all three dimensions, let alone their interactions [Albrecht:2018:00]. As the Intergovernmental Panel on Climate Change (IPCC) 2023 synthesis report states: “Maladaptation can be avoided by flexible, multi-sectoral, inclusive, long-term planning…” [ipcc:2023:00]. Without common ontologies and standards, attempts at integration often result in incoherent or incompatible models [holzinger:2005:00, razavi:2025:00].
Pitfall 3: Bottom-up modeling misses cascading interdependencies. Most disciplines model discipline-specific systems in isolation and only later attempt to couple models. This bottom-up approach, while tractable, is inadequate for understanding cascading effects among systems such as climate, agriculture, energy, land-use, and governance [creutzig:2012:00]. These systems are dynamically coupled, including simultaneous processes and constraints. For example, regional drought in the western U.S. simultaneously constrains hydropower output, increases reliance on natural-gas generation, and raises greenhouse emissions, a cross-sector cascade that purely sectoral models routinely miss [bartos:2015:00]. The Australian Network Infrastructure for Energy, Water, and Hydrogen [Farid:2023:ISC-AP88] project revealed dozens of uncharacterized, high-leverage interactions among infrastructure layers, suggesting the need for coordinated top-down system-wide design merged with bottom-up process-level modeling. When scaled to ten sectors each with 100 subsystems, the potential for cross-sector interactions exceeds one million. Even if only 1% are relevant, thousands of studies would be needed to identify, characterize, and model them [Bi:2022:00, Farid:2022:01]. Fragmented modeling efforts are not only inefficient but blind to emergent risk.
Pitfall 4: Co-simulation limitations in structure and scale. Co-simulation is a popular technique for connecting models of different systems, but it usually assumes a sequential or looped flow of information (i.e., input/output pairs)[gomes:2017:00]. This structure is ill-suited for modeling multilateral, bi-directional SoS with complex feedbacks and distributed agency [taveres-Cachat:2021:01]. Moreover, co-simulation methods are rarely equipped to manage the emergence of disparate spatial and temporal scales that arise in climate systems, infrastructure transitions, and social dynamics [barbierato:2022:00, alfalouji:2023:00]. For example, co-simulating urban energy, water, and transportation networks has proven difficult because decisions about land-use or infrastructure investment unfold over decades, whereas operational control models run on hourly timescales [turner:2017:00, van-dam:2012:00]. In the context of accelerating global change, we need tools that allow for multi-timescale emergence, with the ability to represent slow-evolving institutional transformations alongside rapid technological shocks [lemke:2000:00]. Without this, convergence modeling collapses under either oversimplification or unmanageable complexity.
Pitfall 5: Overlooking socio-psychological and institutional dynamics. Finally, many technical approaches fail to engage the human dimensions of system transformation [van-bruggen:2019:00, sloan:2013:00]. Social license, trust, public values, and cultural narratives are crucial for enabling systemic change, particularly for controversial or disruptive transitions [loorbach:2017:00, wittmayer:2019:00]. For instance, the backlash against wind-farms in rural Europe and the U.S., which is driven more by perceptions of fairness and procedural exclusion than by technology performance, illustrates how neglecting social legitimacy can derail technically sound projects [devine-wright:2011:00, wolsink:2018:00]. Yet these factors are often omitted from modeling frameworks, seen as “soft” or outside the purview of technical experts. The absence of participatory modeling and community co-design leads to brittle solutions that lack legitimacy or adoption [cash:2003:01, kunseler:2015:00, bandari:2024:00, szetey:2023:00]. This dynamic reflects a broader “technical debt” that afflicts the analytical and institutional infrastructure of many sectors [Letovzey:2016:00, Besker:2018:00, Avgeriou:2020:00, Banker:2021:00, Dooley:2023:00]. As evidenced by recent systemic failures (e.g., Southwest Airlines’ operational meltdown), legacy systems often collapse not because of technical flaws, but because they were never designed for systemic adaptability.
Pitfall 6: Not communicating the benefits of convergence Collectively, these five pitfalls reveal not isolated shortcomings, but a shared structural deficiency in the way science organizes knowledge and practice. They expose the absence of a unifying scientific and institutional infrastructure capable of linking discipline-specific tools, data, and governance systems into a coherent whole. Without such integration, existing approaches will continue to generate fragmented, mismatched, or non-transferable solutions, mirroring policy silos in which government agencies pursue parallel mandates while overlooking critical cross-sector dependencies [Cumming:2013:00, morrison:2017:01].
An actionable convergence paradigm must therefore be use-inspired, grounded in shared ontologies, and extensible across scales, systems, and regions, enabling a shared architecture of methods, models, and educational tools co-developed with stakeholders. The alternative is untenable: isolated frameworks, region-bound models, and incremental system coupling cannot meet the urgency or complexity of 21st-century Anthropocene challenges [Bi:2022:00, Farid:2022:01].
3 Meta-Cognition Map as the Scaffold for a Convergent SoS Paradigm
This paper advances an actionable SoS convergence paradigm designed to overcome the limitations of single-discipline and single-problem approaches [Bi:2022:00, Iwanaga:2021:00, Little:2019:00, Little:2016:00]. At the foundation of this paradigm is the structured meta-cognition map (Fig. 3), which serves as a scaffold for knowledge generation across scientific and engineering domains. The map provides a systematic framework for understanding how knowledge is created, translated, and applied across multiple representational layers. It also establishes criteria for evaluating convergence potential, guides the development of methodologies capable of supporting coherent, multi-domain SoS analysis, and informs how to train future adopters of this framework.
Figure 3: Meta-cognition map of scientific knowledge generation, illustrating how individuals think while moving among five complementary forms of cognition: Real-World, Systems Thinking, Visual, Mathematical, and Computational. Adapted from [Farid:2023:01, Little:2025:ISC-JR10].
3.1 Knowledge Generation Described by a Meta-Cognition Map
The meta-cognition map (Fig 3) conceptualizes how individuals and teams generate scientific knowledge across five interdependent domains: Real-World, Systems Thinking, Visual, Mathematics, and Computing. Each of the domains depicted in the meta-cognition map individually generate knowledge that is essential to overall system understanding but limited when isolated. By explicitly linking these domains through common ontological constructs, the meta-cognition map supports transdisciplinary, scalable, and actionable knowledge generation, a necessity for modeling and managing systems of Anthropocene systems.
The Real-World Domain. Starting from the left side of the meta-cognition map (Fig. 3), scientific inquiry begins in the Real-World domain, through physical observation, sensing, and experimentation [flick:2006:00, popper:2005:00]. In many engineering disciplines, especially those dealing with infrastructure or environmental systems, controlled experimentation is often infeasible, prohibitively expensive, or ethically inappropriate. For instance, scientists cannot conduct controlled experiments on the Earth’s climate or river networks as they would in a laboratory setting [clark:2003:00, oreskes:1994:00]. Instead, researchers rely on field measurements, historical records, and increasingly, satellite and remote-sensing technologies [council:2007:00, reichstein:2019:00]. Yet these observational data are often incomplete, noisy, or disconnected in time and space, as is evident in global hydrological or atmospheric monitoring [montanari:2013:00, beven:2018:00]. These limitations necessitate the transition to other domains to extract deeper insight.
The Systems Thinking Domain. When direct experimentation is limited, systems thinking enables researchers to synthesize observations into holistic mental models that reveal interconnections, feedbacks, and emergent behavior [ackoff:1994:00, checkland:2000:00, senge:2006:00]. Within this domain, physical intuition, design thinking, and thought experiments thrive as tools for managing cognitive complexity [espejo:1994:00, cabrera:2008:00]. Yet systems thinking remains largely interpretive and subjective; conceptualizations often differ between individuals and disciplines [midgley:2000:00]. While it is essential for abstraction and creative synthesis, its lack of formal structure limits reproducibility and analytic rigor. To move from intuitive understanding to actionable modeling, systems thinking must be expressed in a structured visual or mathematical language [forrester:1990:00].
The Visual Domain. The visual domain plays a pivotal role in bridging informal thought and formal representation [jankun-kelly:2007:00]. Every discipline has developed its own visual abstractions, including free-body diagrams in mechanics, circuit diagrams in electrical engineering, process flow diagrams in chemical engineering, and GIS maps in geospatial disciplines. These tools enable communication, iteration, and collaborative model building [valencia-garcia:2021:00, al-kodmany:2001:00]. However, they are deeply discipline-specific. This specialization can become a barrier when collaborating across domains, particularly in SoS contexts. Cross-disciplinary collaboration demands that visual tools be ontologically coherent and extensible across systems [aagaardhansen:2007:00]. SysML, for instance, offers a structured visual syntax that can span mechanical, cyber, and environmental systems [mann:2009:00].
The Mathematics Domain. Mathematics provides the rigor needed to move from conceptual to analytical understanding [banerjee:2021:00, gershenfeld:1999:00]. Algebraic formulations, differential equations, optimization models, and statistical representations enable theoretical reasoning and system-level inference across diverse scientific domains [beltrami:2014:00]. Importantly, mathematics allows generalization across classes of systems rather than specific instances. Yet mathematical abstraction also introduces trade-offs: to achieve tractability, models often simplify nonlinearities, ignore uncertainty, or omit contextual detail [box:1979:00, beven:2018:00, saltelli:2020:00]. Moreover, mathematical formalism can fail to capture the full semantics of real-world phenomena unless grounded in a clear ontological foundation [rosen:2011:00, cartwright:1983:00, Farid:2022:ISC-J49]. Thus, it must be translated both forward from conceptualizations and backward into computational models for meaningful application.
The Computing Domain. The computing domain enables researchers to operationalize mathematical models and extend them to large-scale, high-dimensional problems [winsberg:2019:00, hey:2009:00]. Simulation, optimization, and machine learning algorithms form the computational core of modern systems modeling [stummer:2021:00, sayama:2015:00]. High-performance computing platforms allow scientists to test alternative scenarios, identify emergent behaviors, and explore uncertainty across scales. However, computation is not epistemically neutral. Outcomes are highly sensitive to upstream conceptual and mathematical assumptions [oreskes:1994:00, saltelli:2020:00]. Moreover, coding and interpreting complex simulations demand specialized skills that are cognitively distant from natural-language or visual reasoning [miller:2001:00]. For these reasons, the computing domain must remain tightly integrated with the other four domains to ensure transparency, interpretability, and scientific coherence.
The cyclical flow of knowledge generation is captured in the meta-cognition map. Collectively, these five domains form the scaffolding of this framework (Fig. 3). Together, they describe how knowledge is created, translated, and applied, from real-world observation to computational implementation and back again. In a convergence paradigm, inquiry does not progress linearly within a single domain; often, because it runs into the knowledge barrier associated with that domain. Rather, it iterates across all five, moving between them as insights are refined, tested, and re-expressed in new forms. Modeling a complex socio-environmental system thus becomes a process of translation, moving ideas from one representational form to another, each grounded in a shared set of ontological constructs [guizzardi:2006:00, Guizzardi:2005:00]. Conceptualization links the real-world and systems-thinking domains; formalization bridges the visual and the mathematical domains; and implementation occurs within the computing domain.
Once computation produces results, generated knowledge traverses the domains in reverse to produce insights (Fig. 3). Computational outputs are interpreted within the mathematical domain to produce visualized findings through graphs, maps, or schematic representations. Within the systems thinking domain, these visual insights can re-inform mental models of the overall system. Finally, this synthesized understanding returns to the real-world domain, where it informs empirical observation, stakeholder communication, and policy intervention.
Ontological coherence–the preservation of meaning across disciplinary boundaries–is the foundation of this convergence paradigm. While disciplinary models can remain internally consistent when confined to a single domain [nielsen:2015:00], Anthropocene systems inherently span environmental, economic, social, and technological dimensions. Addressing such systems therefore requires a framework that maintains rigor within each domain while enabling consistent translation across them.
Over the past decade, this convergence paradigm, implemented through MBSE and HFGT, has been applied across a wide range of systems. Within individual domains, HFGT has supported analyses of electric power networks [Farid:2015:SPG-J17, Thompson:2021:SPG-J44], drinking water distribution systems [Farid:2014:IEM-C41], transportation infrastructures [Viswanath:2013:ETS-J08], and mass-customized manufacturing and production systems [Farid:2008:IEM-J04, Farid:2008:IEM-J05, Farid:2017:IEM-J13, Farid:2015:IEM-J23].
Its use has also expanded to multi-domain systems of systems, including multi-modal electrified transportation [Farid:2016:ETS-J27, Farid:2016:ETS-BC05, vanderWardt:2017:ETS-J33], microgrid-integrated production systems [Schoonenberg:2017:IEM-J34], hydrologic and watershed systems [harris:2025:00], personalized and geographically distributed healthcare networks [Khayal:2015:ISC-J20, Khayal:2018:00, Khayal:2021:ISC-J46], hydrogen–natural gas co-infrastructure [Schoonenberg:2022:ISC-J48], the energy–water nexus [Farid:2024:ISC-AP101], and the U.S. multi-modal energy system [Thompson:2024:ISC-J53]. Recent work in healthcare delivery for incarcerated individuals [Satcher:2025:ISC-JR05] underscores that, in many Anthropocene contexts, the central challenge lies less in the mathematical representation itself and more in establishing shared mental models and system-level framing.
3.2 Design Criteria for a Convergent System-of-Systems Paradigm
Based on this understanding of knowledge generation, a convergence paradigm must exhibit six essential characteristics to effectively address interconnected Anthropocene challenges. First, it must be grounded in a strong systems foundation, incorporating principles from systems thinking and systems engineering to capture feedback, emergence, hierarchy, and dynamic behavior [senge:2006:00, forrester:1990:00, checkland:2000:00, nielsen:2015:00, weck:2011:00, Little:2016:00, Iwanaga:2021:00]. This foundation enables integrated treatment of physical and social systems [Little:2019:00, Bi:2022:00]. A common ontology is equally vital, providing a shared set of fundamental system elements (e.g., resources, processes, operands) that ensures consistency across visual, mathematical, and computational representations and enables seamless translation across disciplines [Guizzardi:2005:00, guizzardi:2006:00, mann:2009:00, Farid:2022:01, dori:2016:01]. Extensibility further enables the paradigm to scale across system sizes (e.g., local to regional), span multiple domains (e.g., energy, water, transportation), and accommodate growing structural complexity without reformulation [nielsen:2015:00, Little:2019:00, Iwanaga:2021:00, verburg:2016:00].
In addition to these structural requirements, the paradigm must be both analytical and synthetic. Analytical capability supports rigorous evaluation of existing systems using formal models grounded in mathematics [banerjee:2021:00, beltrami:2014:00, gershenfeld:1999:00, box:1979:00, beven:2018:00]. Synthetic capability complements this by enabling generation of new system configurations through architectural redesign, feedback reconfiguration, or policy experimentation [Farid:2016:00, weck:2011:00, dori:2016:01, Little:2016:00]. Finally, the paradigm must be problem independent. Rather than being tailored to a single societal challenge, it should support diverse use cases without fundamental structural changes [yates:2018:00, Albrecht:2018:00, holzinger:2005:00, coccia:2016:00, razavi:2025:00]. Together, these characteristics make the paradigm robust, reusable, and capable of supporting integrated, actionable knowledge across complex, evolving systems [Bi:2022:00, Little:2019:00].
3.3 Analysis and Synthesis Methods with Convergence Potential
While the meta-cognition map provides a conceptual scaffold for cross-domain integration, the realization of an actionable SoS convergence paradigm depends on modeling frameworks that both reflect and reinforce its structure. Specifically, candidate methods must support iteration across the five cognitive domains–Real-World, Systems Thinking, Visual, Mathematics, and Computing–while also satisfying the six design criteria introduced above: a strong systems foundation, a common ontology, extensibility, analytical and synthetic capabilities, and problem independence [Little:2019:00, Bi:2022:00, Iwanaga:2021:00].
Among existing modeling traditions, four methodologies exhibit particularly high potential for enabling such convergence: Network Science, Data-Driven Artificial Intelligence (AI), MBSE, and HFGT. Each engages multiple domains of the meta-cognition map and addresses specific limitations in traditional disciplinary approaches.
Network Science. Network Science provides a generalizable mathematical and visual language for describing relationships and flows within complex systems [Barabasi:2016:00, newman:2018:00]. Nodes and edges can represent physical infrastructure (e.g., rivers, power lines), social connections, or logical pathways (e.g., information or energy exchange). Network models support analytical evaluation of connectivity, resilience, and modularity, and are highly extensible across domains [weck:2011:00]. They align closely with the visual and mathematical domains of the meta-cognition map and are grounded in systems thinking principles such as feedback and emergence [ackoff:1994:00, senge:2006:00]. However, network models – including multi-layer networks – often suffer from onotological limitations [Kivela:2014:00] that limit their application to practical real-world systems that exhibit high levels of heterogeneity. In contrast, MBSE and HFGT in many cases overcome these ontological limitations and facilitate the analysis of highly heterogeneous systems of systems [Farid:2022:01].
Data-Driven Artificial Intelligence (AI). AI methods, including machine learning, natural language processing, and deep learning, excel at uncovering patterns, correlations, and emergent structures from large, complex, and noisy datasets. They extend the computational and real-world domains of the meta-cognition map, offering tools for data assimilation, prediction, and anomaly detection [reichstein:2019:00]. While AI lacks an inherent systems ontology, it can be guided by structured domain knowledge or integrated with ontologically grounded frameworks [oreskes:1994:00, saltelli:2020:00]. AI is inherently analytical, diagnosing patterns and relationship, and increasingly synthetic, as generative models can propose new designs or interventions. Its problem independence makes it highly versatile for convergence modeling, though care must be taken to prevent “black-box” outcomes that obscure causal understanding [beven:2018:00].
Model-Based Systems Engineering (MBSE). MBSE provides a structured, formal approach to representing system architecture, behavior, and requirements through visual and semantic modeling languages such as SysML [wymore:2018:00, ramos:2012:00, mann:2009:00]. It spans the systems thinking, visual, and mathematical domains of the meta-cognition map and directly supports ontological coherence [Guizzardi:2005:00, guizzardi:2006:00]. MBSE enables clear, modular representations of complex systems and allows traceable translation between conceptual, analytical, and computational models [dori:2016:01]. It has strong synthetic capabilities, supporting design and evaluation of new architectures, and is inherently extensible across domains. However, MBSE alone offers limited analytical capability; its power lies in structuring systems prior to integration with analytical frameworks such as HFGT.
Hetero-functional Graph Theory (HFGT). HFGT provides a mathematically rigorous, ontology-driven framework for modeling systems through their functional capabilities rather than component identities [Farid:2022:01, Farid:2016:00]. Buffers (e.g., storages or states) and elements (e.g., transformations or transports) are linked through operand flows, creating a scalable structure that unites system architecture and behavior. HFGT spans all five cognitive domains: it begins with real-world system elements, formalizes systems abstractions, encodes visual and mathematical representations, and enables large-scale simulation through computing [harris:2025:00, naderi:2025:00]. It also satisfies all six convergence characteristics–strong systems foundation, shared ontology, extensibility, analytical and synthetic capacity, and problem independence–making it particularly suitable for multi-domain integration across environmental, social, and engineered systems.
These four methodologies are not mutually exclusive but complementary. Network Science provides scalable abstraction of interconnectivity; AI extracts knowledge from data-rich environments; MBSE structures interdisciplinary insight into coherent system architectures; and HFGT operationalizes those models with ontological rigor and computational power. Together, they form a coherent modeling toolset that supports iterative translation across cognitive domains and equips the next generation of Anthropocene System Integrators with the means to both understand and intervene in complex socio-environmental systems.
3.4 Training the Next Generation of Anthropocene System Integrators
Addressing the intertwined and evolving challenges of the Anthropocene requires a new kind of practitioner: the Anthropocene System Integrator. This subsection translates the cognitive and methodological foundations of the convergence paradigm into human capabilities. These integrators are not defined by a single discipline or method, but by their ability to traverse domains of knowledge, synthesize insight across systems, and co-create actionable solutions. Their core skillset mirrors the domains of the meta-cognition map–Real-World, Systems Thinking, Visual, Mathematics, and Computing–each anchoring a complementary set of competencies essential for convergence [senge:2006:00, clark:2003:00, midgley:2000:00].
Real-World Domain: Observation, Sensing, and Empirical Validation. Anthropocene System Integrators must begin with the world as it is. Skills in physical experimentation, remote sensing, and empirical data collection ground their models in reality [reichstein:2019:00]. These competencies enable integrators to understand dynamic Earth systems, calibrate models, and validate simulations. Given the scale and ethical constraints of many global challenges, these observational methods serve as the evidence-base for model development.
Systems Thinking Domain: Holistic Conceptualization and Mental Modeling. Complex systems cannot be reduced to isolated parts. Systems thinking equips integrators to recognize multiple forms of formal and functional hierarchy, feedback loops, nonlinearity, emergence, and cross-scale interdependencies [ackoff:1994:00, checkland:2000:00, senge:2006:00]. This mental scaffolding enables them to identify leverage points, anticipate unintended consequences, and frame problems in ways that reflect the interconnected structure of the Anthropocene. Most of all, systems thinking is the primary tool for Anthropocene Systems Integrators to understand and ultimately manage complexity.
Visual Domain: Representation, Communication, and Architecture. To transition from intuition to formal design, integrators must represent systems visually. Skills in diagramming, GIS mapping, and system architecture, especially through tools like SysML, allow them to communicate structure and behavior across disciplines [mann:2009:00, dori:2016:01]. Visualization is not only a design tool; it fosters shared understanding among collaborators, bridging the gap between technical expertise and stakeholder intuition.
Mathematics and Computing Domains: Analysis, Simulation, and Scale. The analytical core of an integrator’s toolset lies in mathematical modeling and computation. These skills support the translation of visual architectures into formal equations, algorithms, and simulations. From optimization to machine learning, integrators must be fluent in tools that scale insight across dimensions and timeframes [banerjee:2021:00, winsberg:2019:00, saltelli:2020:00]. They must also understand the assumptions, limitations, and implications of computational methods to avoid “garbage-in-garbage-out” results.
Cross-cutting skills: Translation, Engagement, and Storytelling. Beyond domain-specific capabilities, Anthropocene System Integrators require skills that bridge boundaries. Linguistic, conceptual, and ontological translation skills enable them to adapt models across domains, making technical work legible to collaborators from other fields. Stakeholder engagement is equally essential: integrators must co-create models, elicit values, and build trust through participatory processes [ulibarri:2018:00, falconi:2017:00]. Storytelling complements these efforts by making complexity human. Whether through narratives, metaphors, or visuals, integrators must convey insights in ways that resonate with diverse audiences [cash:2003:01, kunseler:2015:00, loorbach:2017:00]. Translation skills foster flexibility across disciplinary jargon, enabling integrators to reveal conceptual parallels among approaches such as the convergence paradigm, process integration, industrial symbiosis, heating network synthesis, electricity sector coupling, and multi-sector coupling—ultimately helping to overcome siloing caused by domain-specific languages [schipfer:2026:00]. Above all, Anthropocene system integrators must champion systemic synergies while candidly exposing the risks and unintended consequences of growing convergence. These cross-cutting skills are essential for navigating the increasing complexity of systems of systems in a world often polarized between progressive, yet sometimes overly enthusiastic, change makers and conservative, though equally relevant, critics of complexity.
These skills do not stand alone. They form an iterative, reflexive practice of convergence. Integrators observe, conceptualize, model, simulate, interpret, and return to observation, each step informed by a shared ontology and transdisciplinary intent. This feedback loop closes the gap between knowledge generation and action, between abstract analysis and lived experience. It is through this continual translation that convergence becomes operational, making the education of Anthropocene System Integrators central to addressing modern challenges.
4 Case Study: Demonstration of the SoS Convergence Paradigm in the Chesapeake Bay Watershed
(a) Land use
(b) Connected watershed [harris:2026:00]
Figure 4: Chesapeake Bay Watershed
The Chesapeake Bay Watershed serves as a real-world demonstration of how the SoS Convergence Paradigm can be operationalized to real-world Anthropocene challenges. As a coupled human–natural system, the Chesapeake Bay Watershed integrates interacting domains of land use, hydrology, estuarine ecology, economics, and multi-level governance across spatial and temporal scales [he:2019:00, pan:2021:00, phillips:2016:00]. The watershed spans six states and the District of Columbia, channels flows through more than 100,000 tributaries into the nation’s largest estuary, and supports over 17 million residents [pan:2021:00, phillips:2016:00]. Its restoration is coordinated through the Chesapeake Bay Program (CBP), a multi-institutional partnership among federal and state agencies, NGOs, and academic organizations, to address persistent challenges such as hypoxia, eutrophication, and biodiversity loss [Hood:2021:00, boesch:2001:00, webster:2022:00, batiuk:2025:00]. These challenges underscore the limits of sectoral tools and fragmented governance.
From a convergence perspective, the Chesapeake Bay Watershed concentrates the hallmark challenges of Anthropocene systems: scientific complexity, institutional fragmentation, and socio-economic coupling [morrison:2017:01, Hood:2021:00]. The CBP’s modeling system and the Chesapeake Assessment Scenario Tool (CAST) provide valuable data and calibration, yet remain structurally opaque and difficult to extend beyond domain boundaries [Hood:2021:00]. In the CBP context, opacity manifests in routine yet undocumented practices like unofficial spreadsheet exchanges, previewed model runs and negotiated resubmissions, use of state/consultant alternatives (e.g., VAST/MAST), and consensus-based deferral of QA/QC which the dual-layer overlay makes explicit [jangjoo:2025:00]. These conditions call for a framework that satisfies the six design criteria of convergence–systems foundation, common ontology, extensibility, analytical and synthetic capacity, and problem independence–while facilitating translation across the five cognitive domains of the meta-cognition map: Real-World, Systems Thinking, Visual, Mathematics, and Computing.
In the sections that follow, this case study demonstrates how the Convergent Anthropocene Systems (Convergent Anthems) team operationalized the convergence paradigm through four implemented components: 1. an SoS Computational Framework that integrates MBSE, SysML, and HFGT to synthesize, coherently integrate, and validate models of land-use, watershed, and economic systems; 2. an SoS Decision-Support System that combines institutional analysis with MBSE, SysML, and HFGT to ensure that diverse Chesapeake Bay Program stakeholders can meaningfully engage with the computational framework and use it to inform decision-making across regional and urban scales; 3. an SoS Team Science and Collaboration effort that applies principles of Team Science and an evolving Collaboration Agreement to evaluate, document, and facilitate convergence within the interdisciplinary research team; and 4. an SoS Pedagogy that trains the next generation of Anthropocene System Integrators to navigate these domains with fluency, rigor, and collaborative competence.
4.1 SoS Computational Framework
The SoS Computational Framework represents an initial effort toward operationalizing the convergence paradigm. It combines SysML and HFGT to create an open, ontology-grounded environment that unifies the structure and function of hydrologic, economic, and institutional systems. These two methods, identified earlier as possessing high convergence potential, serve as the Visual and Mathematical anchors of the meta-cognition map, translating conceptual structure into executable computation while preserving meaning across representational forms.
Building on prior methodological work, Harris et al. [harris:2025:00] developed a series of simplified watershed reference architectures using SysML and instantiated them in HFGT through case studies ranging from single-lake to three-lake systems. This work established the ontological and mathematical foundations for representing hydrological processes as hetero-functional networks, in which system elements (resources and operands) interact through transformation and transportation relations. These relations include the application of nitrogen via fertilizer and the flow of water and nitrogen from land, to lakes, and through rivers.
This approach was then extended to the full Chesapeake Bay Watershed (Fig. 4). A SysML–HFGT instantiation now captures the hierarchical structure of elements of the watershed, including land segments, outlet points, river networks, and estuarine boundaries [harris:2026:00]. Geospatial features from CAST were translated into tensor-based HFGT representations that were then used in coordination with CAST behavioral data to simulate the movement of nitrogen and phosphorus throughout the watershed (Fig. 4(b)). The project utilizes the Weighted Least Squares Error Hetero-functional Graph State Estimator (WLSE–HFGSE) introduced by Thompson et al. [Thompson:2025:ISC-JR11] to overcome the uncertainty inherent in large, data-sparse systems such as in the CAST model . When adapted to a watershed system, this estimator infers unobserved water and nutrient flows from partial data while maintaining physical consistency [harris:2026:00]. The result was a transparent, extensible model that mirrors CAST’s routing assumptions while remaining fully aligned with MBSE principles. Rather than replicating CAST, this ongoing modeling effort seeks to augment it, revealing structural relationships withing the watershed system and establishing a foundation for future extension to additional environmental or socio-technical systems.
In parallel, Naderi et al. [naderi:2026:00] demonstrated the methodological generality of the SysML–HFGT framework by applying it to economic input–output (I–O) models. Using a Rectangular Choice of Technology (RCOT) example [Duchin:2011:00], they showed that conventional I–O dynamics can be reproduced within HFGT’s mathematical formalism, providing structural clarity and enabling integration with environmental systems. This offers an initial application of the convergence framework to economic process modeling for future extension to the Chesapeake Bay Watershed region to coordinate with hydrological processes.
Many computational approaches for modeling complex systems of systems, such as those described in this case study, have traditionally relied on system dynamics. Naderi et al. [naderi:2025:00] examined how the integrated HFGT-MBSE framework compliments and differs from system dynamics. While traditional system dynamics models have long been central to human–natural systems analysis, they employ a limited set of systems abstractions that can constrain cross-domain scalability. SysML and HFGT, in contrast, (a) encode a broader spectrum of systems-thinking constructs, (b) reproduce analytical conclusions obtained via system dynamics, and (c) overcome methodological barriers to multi-domain integration [naderi:2025:00]. SysML offers the graphical ontology and architectural traceability; HFGT provides the mathematical ontology and computational rigor. Together, they form the convergence spine that links observation to simulation, structure to function, and analysis to design.
Through the lens of the meta-cognition map, this framework traverses the Visual domain (SysML diagrams), Mathematical domain (HFGT tensor calculus), and Computing domain (simulation and estimation), all grounded in Real-World data and guided by Systems Thinking principles. Ontological coherence ensures that meaning and physical validity are preserved throughout translation [Guizzardi:2005:00, guizzardi:2006:00].
4.2 SoS Decision-Support System
In parallel with the computational framework, the Convergent Anthems team developed an SoS Decision-Support System to connect technical modeling with the institutional and policy dynamics of the Chesapeake Bay restoration effort. Designed for coherence with the SysML–HFGT computational framework, it integrates stakeholder perspectives across regional and urban scales and translates analytical results into actionable governance insights.
Three complementary protocols operationalize this system. The first integrates institutional analysis with SysML to map and diagnose gaps between formal BMP policies (“rules-in-form”) and their practical implementation (“rules-in-use”) [jangjoo:2025:00]. The process formalizes the CBP BMP verification workflow in a SysML activity diagram with actor swimlanes and control flows (rules-in-form), then overlays empirically extracted capability statements structured as actor + verb + object [condition] from interviews, observations, and trace documents (rules-in-use). This research defines empirically observed capability paths (rules-in-use) that bypass, run in parallel to, or repurpose the prescribed SysML activity flow (rules-in-form) as divergence nodes. These divergent nodes are classified as drift, layering, conversion, or displacement [jangjoo:2025:00]. This analysis shows how institutional drift, layering, conversion, and displacement are enacted through socio-technical artifacts that systematically favor well-resourced actors and reproduce capacity disparities. However, rather than framing divergence as mere implementation failure, the study demonstrates that informal adaptations often represent rational strategies to manage resource constraints and political risks. Technical routines, far from being neutral, operate as boundary objects and conduits of hidden influence, shaping governance outcomes in subtle but consequential ways.
The second effort, the Analysis of Topic Model Networks (ANTMN) [walter:2019:00], applies machine learning and network analysis to study how conversations within the Chesapeake Bay Program have evolved over the past two decades. The approach analyzes 29,177 CBP committee documents (2004–2024) to identify recurring themes in stakeholder discourse and track how policy priorities are reallocated over time. To do this, the team applied a text-mining method known as Latent Dirichlet Allocation (LDA) to inductively identify topics based on word co-occurrence patterns [blei:2003:00]. These topics were constructed into a topic network to capture how topics appeared together in the same documents, and community detection was applied to group related topics into broader themes. Findings indicate that adaptive governance in the CBP does not steadily grow through learning and action as often assumed. When technical changes—especially updates to the CBP’s modeling tools—occur, they dominate the agenda and discourse, pushing out discussion of the organization’s broader vision and community engagement. These model-driven shifts had a stronger impact than external events such as presidential transitions and the COVID-19. When the results were summarized into broader themes, clear patterns emerged: CAST Phases 5.3 and 5.3.2 produced the largest and most persistent reallocations of attention, where conversations shifted from long-term visioning toward practical implementation. Accountability checks tend to trail implementation, making oversight reactive rather than proactive. Overall, findings suggest that managing attention—alongside resources—is necessary to keep governance balanced and prevent it from becoming overly technical at the expense of an organization’s original mission.
The third effort extends the decision-support system to the coordination of the enterprise responsible for implementing restoration actions. In practice, the Chesapeake Bay Program operates as a large, multi-agency socio-environmental enterprise whose scale, cross-jurisdictional dependencies, and resource constraints resemble those of an engineering mega-project [Hosseini:2025:ISC-J55]. At this level, effective decision support must account not only for real-world system behavior but also for the architecture, scheduling, and resource allocation of the implementing organization itself. However, conventional resource-constrained project scheduling (RCPSP) approaches typically neglect explicit representation of enterprise structure, contributing to the well-documented “iron law of mega-projects,” in which initiatives frequently exceed cost and schedule targets [Hosseini:2025:ISC-J55, Hosseini:2025:ISC-JR17]. To address this limitation within the same ontological framework used for hydrologic and institutional modeling, the SoS convergence paradigm applies MBSE and HFGT through a hetero-functional network minimum-cost flow formulation that generalizes RCPSP. This enables enterprise-level coordination to be analyzed using the same formalism as physical and governance processes, linking system dynamics to actionable program management.
To ensure validity, the institutional layer underwent member checking with CBP practitioners (diagram-review workshops), independent double-coding of interview materials with line-by-line reconciliation, and digital trace tie-back linking each informal capability to time-stamped artefacts. These procedures create an auditable chain between narrative evidence and model elements and align with best practices for transparent process-tracing [jangjoo:2025:00].
Together, these protocols link institutional structure, social narrative, and computational modeling within a common ontology, advancing the analytical, synthetic, and problem-independent capacities of convergence. Viewed through the meta-cognition map, the Decision-Support System spans the Real-World, Systems Thinking, and Computing domains, strengthening the bridge between scientific modeling and stakeholder decision-making.
4.3 SoS Team Science and Collaboration
Effective convergence requires not only shared technical frameworks but also deliberate cultivation of team processes that sustain integration across disciplines, institutions, and geographies. To this end, the Convergent Anthems project employed a structured Team Science approach to evaluate and facilitate collaboration within the interdisciplinary team. This approach provided the social infrastructure necessary for implementing the SoS Convergence Paradigm and training the next generation of Anthropocene System Integrators.
Building on the principles of the Toolbox Dialogue Initiative and the framework developed by Bennett et al. [bennett:2018:00], the team co-developed a Collaboration Agreement that formalized expectations for communication, authorship, and team governance. The agreement emerged through facilitated workshops, iterative surveys, and consensus-based discussion during multiple “Team of Teams” meetings. It was designed to be a living document–revisited and refined annually–to facilitate and maintain transparency, trust, and adaptability as the team evolved. To assess progress toward convergence, the team conducted periodic surveys measuring key dimensions of team coherence including team member perceptions of conflict, trust, and cohesion, following established constructs in the Team Science literature [fiore:2015:00].
Viewed through the meta-cognition map, these Team Science practices support translation across all five domains. They enable shared understanding in the Real-World domain through collaborative governance, cultivate collective Systems Thinking, standardize representation in the Visual and Mathematical domains via SysML and HFGT, and coordinate efforts within the Computing domain. This alignment promotes the progression of social and technical integration in partnership.
4.4 SoS Pedagogy
The SoS Educational Pedagogy was developed to translate the Convergence Paradigm into a scalable training model that cultivates the next generation of Anthropocene System Integrators. These integrators must not only master technical modeling tools but also develop the reflexive and collaborative skills necessary to operate across the five cognitive domains of the meta-cognition map. The SoS Pedagogy therefore combines technical instruction, systems thinking, and team-based learning within a framework that is both analytically rigorous and transdisciplinary in practice.
Graduate students and early-career researchers continue to be trained in MBSE using SysML and HFGT, with assignments tailored to their focus on computational modeling or institutional analysis. Training modules were deployed across multiple institutions, including Virginia Tech and Stevens Institute of Technology, to ensure pedagogical consistency while maintaining disciplinary relevance. At Virginia Tech, CEE 4134/5114 Sustainable Systems introduced SysML and HFGT through project-based modules on Mono Lake and multi-lake systems, demonstrating the frameworks’ extensibility from simple to complex watershed configurations [naderi:2025:00, harris:2025:00]. At Stevens, SE 650 Systems Architecture and Design centered on the meta-cognition map as the conceptual foundation for systems modeling, with student teams applying SysML to Anthropocene-scale challenges such as transit networks, hydrogen integration, and urban carbon management.
A central resource in this pedagogy is a laboratory manual which serves as a structured guide to convergence research practices encompassing data management, software workflows, model documentation, and research dissemination. Serving as a boundary object for convergence, the manual helps new researchers adopt best practices in reproducible modeling and collaborative system design, reinforcing ontological coherence and computational integrity across teams.
Together, these educational efforts have established a replicable model for convergence training that links conceptual understanding to computational implementation. Students learn to move iteratively through the five cognitive domains, observing and defining problems in the Real-World domain, framing them through Systems Thinking, formalizing them visually and mathematically in SysML and HFGT, and finally operationalizing them in the Computing domain through simulation and estimation. This domain traversal is explicitly reinforced through project-based learning and collaborative reflection, equipping trainees to both build and interpret complex models.
Viewed within the broader SoS Convergence Paradigm, the SoS Pedagogy functions as the educational backbone of the framework. It not only equips researchers with modeling fluency but also fosters metacognitive awareness, shared language, and transdisciplinary empathy. Through this pedagogy, the convergence paradigm becomes not merely a research framework but a transferable educational model that can be replicated across institutions and domains.
4.5 The Chesapeake Bay Watershed as a Convergence Testbed
Together, these elements demonstrate the feasibility of implementing the SoS Convergence Paradigm in a real-world system-of-systems context. The SoS Computational Framework, SoS Decision-Support System, and SoS Pedagogy collectively embody the six design characteristics of convergence while operationalizing the iterative, multi-domain process described by the meta-cognition map (Fig. 3).
A key enabler of this integration has been SysML, now used ubiquitously across the Convergent Anthems project. As researchers from hydrology, economics, governance, and data science converge on shared problems, SysML provides a common graphical language for reconciling diverse theories, data structures, and disciplinary vocabularies. Its diagrams act as boundary objects that externalize mental models, expose ontological differences, and facilitate transparent dialogue among collaborators. Within this shared visual domain, the meaning of each system element, whether a nutrient buffer, policy actor, or data process, can be discussed, refined, and ultimately grounded in a consistent ontology compatible with HFGT. SysML thus links the qualitative reasoning of human systems with the quantitative rigor of computational modeling, bridging the visual and mathematical domains of the meta-cognition map.
As demonstrated in prior work, SysML and HFGT together provide the methodological “spine” of convergence modeling: SysML formalizes conceptual and structural relationships through its graphical syntax, while HFGT extends these representations into analyzable and computable mathematical form. This dual framework enables interdisciplinary teams to develop shared visual mental models, a feature shown to be essential for solving complex socio-technical problems [andrews:2023:00].
Through these integrated efforts, the Chesapeake Bay Watershed becomes more than a regional case but rather an exemplar of how convergence science can move from theory to practice. Its combination of structural modeling, institutional analysis, and collaborative pedagogy demonstrates how diverse expertise can be aligned within a coherent ontological and computational architecture. This implementation establishes a transferable template that can be applied to other regions and domains.
5 Conclusions
This paper introduced a novel, actionable SoS Convergence Paradigm for addressing the complex, interdependent challenges of the Anthropocene. Grounded in a meta-cognition map and operationalized through SysML and HFGT, the paradigm provides a unifying framework for cross-domain integration across environmental, economic, and governance systems.
Through the Chesapeake Bay Watershed case study, this paper demonstrated how the paradigm enables structural and functional coherence across heterogeneous models, supports stakeholder-informed decision-making, and bridges the gap between disciplinary specialization and system-level insight. The extensibility of the framework has facilitated its application beyond the Chesapeake Bay, informing initiatives such as the National Energy Analysis Centre (NEAC) at CSIRO [csiro:2025:00] and applications in megaproject planning [Hosseini:2025:ISC-J55].
Because the framework is grounded in a shared SysML–HFGT ontology, it readily accommodates additional subsystems and scales within the Chesapeake Bay Watershed region, including estuarine, governance, and urban contexts. The same computational and decision-support components demonstrated here can therefore be extended from watershed to city and from physical to institutional processes without structural redesign. These properties highlight the paradigm’s generality beyond this case study and position it as a transferable template for convergence-based modeling of other regional and socio-environmental systems.
The central contribution of this work is the demonstration of the convergence paradigm as an extensible modeling architecture for diverse regions and systems. Collectively, the technical, institutional, and educational demonstrations of the paradigm provide a replicable model for convergence science, one that couples the analytical rigor of engineering with the reflexive capacity of systems thinking. By aligning the tools, languages, and teams required for integration, the SoS Convergence Paradigm offers a scalable pathway for advancing resilience, equity, and sustainability in an increasingly interconnected world.
Open Research Section
No new data were created or analyzed in this study.
Conflict of Interest declaration
The authors declare there are no conflicts of interest for this manuscript.
Acknowledgements.
This research is based on work supported by the Growing Convergence Research Program of the National Science Foundation under Grant Numbers OIA 2317874 and OIA 2317877.